# Chiara Amorino, Vytautė Pilipauskaitė

### v2

@@ Abstract
Keywords: Kinetic interacting particle system, parameter estimation, hypoelliptic diffusion, McKean-Vlasov equation, asymptotic normality
@@ 1 Introduction
For $N$ approaching infinity, the interacting particle system described above naturally converges to its mean field equation, represented by the McKean-Vlasov SDE:
@@
where $(\mu_{0},\sigma_{0})$ are the multidimensional parameters we aim to estimate. We refer to $Y^{i}_{t}$ and $X^{i}_{t}$ as the smooth and rough coordinates, respectively. In the absence of interaction, diffusion processes of the form (3) are commonly known as integrated diffusions (see, for instance, [27] or [35]). A notable instance of such models arises when the drift is given by $b(y,x)=-c(y,x)x-\nabla U(y)$ for some function $c$ and potential $U$, and the diffusion coefficient is constant. In this setting, the model (3) is known as a stochastic damping Hamiltonian system, describing the motion of a particle under the combined influence of potential, dissipative, and random forces (see, e.g., [84]). A classical example, in the absence of interaction, is the Kramers oscillator. Among hypoelliptic interacting diffusions, the most relevant example arising from applications and satisfying our hypothesis is the interacting underdamped Langevin dynamics, describing a system of $N$ particles evolving in a confining potential, coupled through an interaction potential, and subject to both noise and dissipation (see Chapter 6 of [70]).
@@ 1.1 Previous literature
A large number of contribution is concerned with statistical inference for diffusions, we refer for example to the books [45, 55] and [58]. Moving in particular to parameter estimation for SDEs, a natural approach to estimating unknown parameters from the continuous observation of the diffusion would be to use maximum likelihood estimation. However, the likelihood function based on the discrete sample is not tractable, as it depends on the transition densities of the process, which are not explicitly known. To overcome this difficulty, several methods have been developed for high-frequency estimation of discretely observed classical SDEs. A widely-used method involves considering a pseudo- likelihood function, often based on the high-frequency approximation of the process dynamics using the Euler scheme, as seen in [29, 54] and [85].
@@
Significant contributions to nonparametric or semiparametric statistical inference in this context can be found in works such as [3, 7, 24], and [67]. Meanwhile, the problem of parameter estimation from observations of both the interacting particle system and the associated McKean-Vlasov equation has been explored in various frameworks and asymptotic regimes; see, for instance, [4, 8, 31, 32, 33, 61, 77, 83] and the references therein.
@@
Melnykova [63] later introduced a contrast estimator based on local linearization (LL), compared it with the least squares estimator, and established a joint central limit theorem. More recently, Gloter and Yoshida [37, 38] introduced adaptive and non-adaptive methods for hypoelliptic models, proving asymptotic normality for estimators based on complete observations. Their local Gaussian (LG) estimator generalizes the LL approach and builds upon [26], but with fewer restrictions on the model class. They also demonstrated that using the full SDE in the contrast function reduces the asymptotic variance of the diffusion estimator, compared to methods using only the rough coordinate.
@@ 1.2 Our contribution
It is important to note that the presence of interaction introduces numerous novel aspects and challenges. Firstly, our asymptotic regime differs from the classical regime commonly found in the literature on classical diffusion. In our case, the time horizon $T$ is fixed. Nonetheless, we are still able to consistently estimate the drift and diffusion coefficients due to the increasing number of particles under consideration. In particular, the number of particles $N$ assumes a role analogous to that played by $T$ in classical diffusion references, and similar conditions on the discretization step emerge (though these now involve $N$ instead of $T$; see Remarks 3.6 and 3.7 for details). Consequently, our approach does not require the existence of an invariant measure to obtain our main results. Our asymptotic results are derived under the condition $N\rightarrow\infty$, crucially leveraging the propagation of chaos.
In this context, we propose two contrast functions tailored to the cases of complete and partial observations, respectively. Due to the non-invertibility of our diffusion matrix, the standard approach of constructing a quasi-likelihood based on the Euler-Maruyama discretization (as in [4]) is not applicable. Nonetheless, our estimation procedure is still anchored in the Euler scheme, though applied differently. Specifically, we focus on estimating the drift and diffusion parameters of the second component of each particle, leading us to propose a contrast function solely based on the Euler approximation of this second component.
In the (most involving) case of partial observations, where only $Y^{i}$ is available, a correction term must be introduced in the contrast function. Our approach, detailed in Section 3, involves replacing the observation of $X^{i}_{t_{j}}$ at each time $t_{j}$ with the increment of the rate process $\tilde{X}^{i}_{t_{j}}:=\frac{1}{\Delta_{n}}(Y^{i}_{t_{j+1}}-Y^{i}_{t_{j}})$, where $\Delta_{n}$ is the discretization step.
@@
The outline of the paper is as follows. In Section 2, we introduce the model and outline the assumptions necessary for our main results, along with an explanation of how the propagation of chaos applies in this context. Section 3 presents our estimators and their properties, which constitute the core of our results. In Section 4, we discuss Hörmander’s condition and explain the transition from model (1) to (3), which is the model used for our estimation procedure. Section 5 introduces several key results that are required for proving our main findings, while Section 6 is dedicated to the proofs of these main results. Lastly, Section 7 focuses on the detailed proofs of all the technical results introduced throughout the paper.
@@ 2 Model and assumptions
We emphasize that, throughout our analysis, the time horizon $T$ will be fixed, and we will consider the asymptotic regime where only the number of particles $N$ tends to infinity.
A key point in our analysis is that we observe discrete observations of the interacting particles, but we can transition from the interacting particle system described above to a system of independent particles through what is known as the propagation of chaos. The system of independent particles consists of $N$ i.i.d. copies of the McKean-Vlasov equation associated with the model (5), as detailed below.
@@
$B(z,\pi):=\begin{pmatrix}x\\ b(z,\pi)\end{pmatrix}\in\mathbb{R}^{2},\qquad A(z,\pi):=\begin{pmatrix}0&0\\ 0&a(z,\pi)\end{pmatrix}\in\mathbb{R}^{2\times 2}$
@@
We recall we assume that these coefficients depend on two parameters, $\mu_{0}$ and $\sigma_{0}$, which we aim to estimate. Therefore, we will denote the drift coefficient as $b_{\mu_{0}}$ and the diffusion coefficient as $a_{\sigma_{0}}$. Furthermore, we will refer to the pair of parameters as $\theta_{0}:=(\mu_{0},\sigma_{0})$, which belongs to the interior of a set $\Theta:=\Theta_{1}\times\Theta_{2}$, where $\Theta_{i}\subset\mathbb{R}^{p_{i}}$, $i=1,2$, are compact and convex.
@@ 3 Estimators and main results
This brings us to the definition of the contrast function within the two observation frameworks under consideration. Recall that, when continuous observation of the process is available, the maximum likelihood estimator performs exceptionally well for parameter estimation. However, when only a discretized version of the process is observed, the transition density (and consequently the likelihood) is not available, making estimation through the maximum likelihood method infeasible.
As pointed out in the introduction, a classical approach to overcoming this issue is to propose a quasi-likelihood function based on the Euler-Maruyama discretization of the process. Even in the case of complete observation, it is not feasible to directly apply the two-dimensional Euler contrast function for each particle $i$ to estimate the parameters $(\mu_{0},\sigma_{0})$. Specifically, the two-dimensional Euler-Maruyama approximation for any $i\in\{1,\dots,N\}$ is given by:
$\begin{pmatrix}Y^{i}_{(j+1)\Delta_{n}}\\ X^{i}_{(j+1)\Delta_{n}}\end{pmatrix}=\begin{pmatrix}Y^{i}_{j\Delta_{n}}\\ X^{i}_{j\Delta_{n}}\end{pmatrix}+\Delta_{n}B(Z^{i}_{j\Delta_{n}},\Pi^{N}_{j\Delta_{n}})+\sqrt{\Delta_{n}}A(Z^{i}_{j\Delta_{n}},\Pi^{N}_{j\Delta_{n}})\begin{pmatrix}\xi^{i}_{j,1}\\ \xi^{i}_{j,2}\end{pmatrix}$,
@@
where $\theta=(\mu,\sigma)$ and the superscript $C$ indicates ”complete observations”. This function is an extension of the classical Euler contrast for unidimensional interacting particle systems, as proposed in [4]. Starting from such contrast function, we define a minimum contrast estimator $\hat{\theta}_{n}^{N,C}$ for complete observations as
@@
where P stands for ”partial observations”. Compared to the contrast function proposed for complete observations, the two key differences are the additional factor of $\frac{3}{2}$, which corrects for the dependence structure highlighted above (and further detailed in Remark 5.5), and a shift in the index of the drift term. It is important to note that $(\tilde{X}^{i}_{j\Delta_{n}})$ is not a Markov process, and the shift in the index of the drift coefficient has been introduced to avoid a correlation term between $(\tilde{X}^{i}_{(j+1)\Delta_{n}}-\tilde{X}^{i}_{j\Delta_{n}})$ and a function of $(\tilde{Z}^{i}_{j\Delta_{n}},\tilde{\Pi}^{N}_{j\Delta_{n}})$. Although this correlation term is small (of the order $\sqrt{\Delta_{n}}$), it is not negligible in our analysis, leading us to the contrast function as defined above.
@@
Assumptions A1-A5 are necessary to establish the consistency of our estimator as detailed in Theorem 3.1. These assumptions are fairly standard in the statistical analysis of random processes. However, the identifiability condition outlined in Assumption A5 warrants particular attention. The quantities $I(\theta)$ and $J(\sigma)$, which are central to this assumption, are not explicitly defined due to their dependence on the underlying law, making this condition challenging to verify in practice. Consequently, ensuring the identifiability of all parameters may not always be feasible.
This difficulty is not unique to our context; it also arises in parameter estimation for interacting particle systems and is not directly linked to the issue of degeneracy in the diffusion coefficient that we address. Section 2.4 of [23] offers an in-depth discussion of this challenge. Specifically, for drift estimation from continuous observations, the authors provide explicit criteria for achieving both identifiability and non-degeneracy of the Fisher information matrix. They establish a connection between global identifiability and the non-degeneracy of the Fisher information for a particular type of likelihood, as highlighted in [23, Proposition 16]. In our framework, this issue remains an open question for further investigation.
@@
(Invertibility) We define a $p\times p$ block diagonal matrix $\Sigma(\theta_{0}){:}=\operatorname{diag}(\Sigma^{(1)}(\theta_{0}),\Sigma^{(2)}(\theta_{0}))$ whose main-diagonal blocks $\Sigma^{(j)}(\theta_{0})=(\Sigma^{(j)}_{kl}(\theta_{0}))$ are defined via
$\Sigma^{(j)}_{kl}(\theta_{0}){:}=\begin{cases}\bar{\Pi}\Big{(}\frac{2\partial_{\mu_{k}}b_{\mu_{0}}\partial_{\mu_{l}}b_{\mu_{0}}}{c_{\sigma_{0}}}\Big{)},\qquad&j=1,\,k,l=1,\dots,p_{1},\\ \bar{\Pi}\Big{(}\frac{\partial_{\sigma_{k}}c_{\sigma_{0}}\partial_{\sigma_{l}}c_{\sigma_{0}}}{c^{2}_{\sigma_{0}}}\Big{)}\qquad&j=2,\,k,l=1,\dots,p_{2}.\end{cases}$
We assume that $\operatorname{det}(\Sigma^{{(j)}}(\theta_{0}))\neq 0$, $j=1,2$.
@@
$(\sqrt{N}(\hat{\mu}_{n}^{N,C}-\mu_{0}),\sqrt{N/{\Delta_{n}}}(\hat{\sigma}_{n}^{N,C}-\sigma_{0}))\xrightarrow{d}\mathcal{N}(0,[\bar{\Pi}\left(\left(\frac{\partial_{\mu}b}{a}\right)^{2}\right)]^{-1})\otimes\mathcal{N}(0,2[\bar{\Pi}\left(\left(\frac{\partial_{\sigma}c}{c}\right)^{2}\right)]^{-1}).$
@@
$(\sqrt{N}(\hat{\mu}_{n}^{N,P}-\mu_{0}),\sqrt{N/{\Delta_{n}}}(\hat{\sigma}_{n}^{N,P}-\sigma_{0}))\xrightarrow{d}\mathcal{N}(0,[\bar{\Pi}\left(\left(\frac{\partial_{\mu}b}{a}\right)^{2}\right)]^{-1})\otimes\mathcal{N}(0,\frac{9}{4}[\bar{\Pi}\left(\left(\frac{\partial_{\sigma}c}{c}\right)^{2}\right)]^{-1}).$
@@
The proof of the main results is provided in Section 6 and centers on analyzing functionals of $Z^{i}_{t_{j}}=(Y_{t_{j}}^{i},X_{t_{j}}^{i})$ and $\tilde{Z}_{t_{j}}=(Y_{t_{j}}^{i},\tilde{X}_{t_{j}}^{i})$ for $i\in\{1,\ldots,N\}$ and $j\in\{1,\ldots,n-1\}$. The asymptotic analysis of these functionals proves to be challenging due to the involvement of both the Markovian process $Z^{i}_{t_{j}}$ and the non-Markovian process $\tilde{Z}^{i}_{t_{j}}$. This analysis will help to understand the asymptotic behavior of the contrast function and its derivatives in both complete and partial observation scenarios, that is the key to obtain our main results. Specifically, we will demonstrate that the first derivatives of the contrast function converge in law to a Gaussian distribution, while the second derivatives converge in probability to a matrix that plays a role in the asymptotic variance of the estimators.
Remark 3.4.
It is important to highlight that when the diffusion matrix is degenerate, as in our model (5), reducing the contrast to the Euler approximation of only the second coordinate still results in the appearance of the law $\bar{\Pi}$ in the variance of the asymptotic Gaussian, rather than the marginal law of the second component as one could have expected.
A parallel can be drawn between our analysis, which involves a growing number of particles $N$ over a fixed time horizon $T$, and the same estimation problem when $N=1$ but $T$ tends to infinity, studied in [54] and [75] (see Remark 3.6 for details).
Notably, when $(X_{t})$ is an autonomous diffusion, as in [54], the asymptotic distribution depends on the stationary distribution of only the diffusion $(X_{t})$. In the context of parameter estimation from a discrete path of a hypoelliptic diffusion over an ergodic time horizon, as studied in [75], the authors observe that the Gaussian variance depends instead on the invariant density of the vector $(X_{t},Y_{t})$.
The parallelism between $N$ and $T$ underscores how our findings align with those presented in [75].
Remark 3.5.
One might wonder whether the estimators we proposed are efficient within the framework under consideration. Thanks to [23], who established the Local Asymptotic Normality (LAN) property for drift estimation in $d$-dimensional McKean-Vlasov models under continuous observations, we know the form of the Fisher information matrix. This form aligns with our results when restricted to drift estimation, confirming the efficiency of our drift estimator.
However, regarding the joint estimation of both the drift and diffusion coefficients in our framework, the LAN property has not yet been proven. Nevertheless, due to the parallelism with the scenario where one estimates parameters for $N=1$ with $T\to\infty$ (see Remark 3.6), one could expect the Fisher information in this case to correspond to the matrix presented in Theorem 3.2 of [4] (see also the discussion on efficiency following Theorem 3.2 therein for more details).
This suggests that, while the estimation of the diffusion coefficient may not be efficient in the case of partial observation—given that its asymptotic variance is inflated by a factor of $\frac{9}{4}$—it could potentially be efficient in the complete observation case. A rigorous analysis of this hypothesis remains a subject for future investigation.

### v3

# Kinetic interacting particle system: parameter estimation from complete and partial discrete observations

@@ Abstract
Keywords: Kinetic interacting particle system, parameter estimation, hypoelliptic diffusion, McKean–Vlasov equation, asymptotic normality
@@ 1 Introduction
For $N$ approaching infinity, the interacting particle system described above naturally converges to its mean field equation, represented by the McKean–Vlasov SDE:
@@
where $(\mu_{0},\sigma_{0})$ are the multidimensional parameters we aim to estimate. We refer to $Y^{i}_{t}$ and $X^{i}_{t}$ as the smooth and rough coordinates, respectively. While the theoretical reduction to (3) is guaranteed by the implicit function theorem, the change of variables is generally not explicit. Consequently, estimating the parameters of the original system given observations of the integrated one (and vice versa) is a highly non-trivial task in practice. For this reason, we proceed by working directly under the assumption that the system takes the explicit integrated form (3). We stress that this is a standard structural assumption in the statistical inference of hypoelliptic diffusions, commonly required even in classical settings without interacting particles (cf. [77, Assumption (C2)]). In the absence of interaction, diffusion processes of the form (3) are commonly known as integrated diffusions (see, for instance, [27] or [37]). A notable instance of such models arises when the drift is given by $b(y,x)=-c(y,x)x-\nabla U(y)$ for some function $c$ and potential $U$, and the diffusion coefficient is constant. In this setting, the model (3) is known as a stochastic damping Hamiltonian system, describing the motion of a particle under the combined influence of potential, dissipative, and random forces (see, e.g., [86]). A classical example, in the absence of interaction, is the Kramers oscillator. Among hypoelliptic interacting diffusions, the most relevant example arising from applications and satisfying our hypothesis is the interacting underdamped Langevin dynamics, describing a system of $N$ particles evolving in a confining potential, coupled through an interaction potential, and subject to both noise and dissipation (see Chapter 6 of [72]).
@@ 1.1 Previous literature
A large number of contribution is concerned with statistical inference for diffusions, we refer for example to the books [47, 57] and [60]. Moving in particular to parameter estimation for SDEs, a natural approach to estimating unknown parameters from the continuous observation of the diffusion would be to use maximum likelihood estimation. However, the likelihood function based on the discrete sample is not tractable, as it depends on the transition densities of the process, which are not explicitly known. To address this challenge, several methods have been developed for the high-frequency estimation of discretely observed classical SDEs. In this context, high-frequency refers to the asymptotic regime in which the discretization step $\Delta_{n}$ tends to zero as the number of observations $n$ increases. A widely-used method involves considering a pseudo-likelihood function, often based on the high-frequency approximation of the process dynamics using the Euler scheme, as seen in [31, 56] and [87].
@@
Significant contributions to nonparametric or semiparametric statistical inference in this context can be found in works such as [3, 7, 24], and [69]. Meanwhile, the problem of parameter estimation from observations of both the interacting particle system and the associated McKean–Vlasov equation has been explored in various frameworks and asymptotic regimes; see, for instance, [4, 8, 33, 34, 35, 63, 79, 85] and the references therein.
@@
Melnykova [65] later introduced a contrast estimator based on local linearization (LL), compared it with the least squares estimator, and established a joint central limit theorem. More recently, Gloter and Yoshida [39, 40] introduced adaptive and non-adaptive methods for hypoelliptic models, proving asymptotic normality for estimators based on complete observations. Their local Gaussian (LG) estimator generalizes the LL approach and builds upon [26], but with fewer restrictions on the model class. They demonstrated that employing a contrast function based only on the rough coordinate increases the asymptotic variance of the diffusion estimator, as the optimal variance is attained only when the full SDE is considered. A similar result was established in [73].
@@ 1.2 Our contribution
It is important to note that the presence of interaction introduces numerous novel aspects and challenges. Firstly, our asymptotic regime differs from the classical framework typically encountered in the SDE literature. In the classical regime, one usually considers a single process ($N=1$) over an increasing time horizon ($T\to\infty$), often requiring the existence of an invariant measure for consistency. In our setting, the time horizon $T$ is fixed. Nonetheless, we are able to consistently estimate the drift and diffusion coefficients due to the increasing number of particles under consideration. Specifically, the number of particles $N$ plays a role analogous to that of $T$ in classical diffusion references, and similar conditions on the discretization step emerge, albeit involving $N$ instead of $T$ (see Remarks 3.4 and 3.5 for details). Consequently, our approach does not rely on ergodicity to obtain the main results. Our asymptotics are derived as $N\to\infty$, crucially leveraging the propagation of chaos.
In this context, we propose two contrast functions tailored to the cases of complete and partial observations, respectively. Due to the non-invertibility of our diffusion matrix, the standard approach of constructing a quasi-likelihood based on the Euler–Maruyama discretization (as in [4]) is not applicable. Nonetheless, our estimation procedure is still anchored in the Euler scheme, though applied differently. Specifically, we focus on estimating the drift and diffusion parameters of the second component of each particle, leading us to propose a contrast function solely based on the Euler approximation of this second component.
In the (most involving) case of partial observations, where only $Y^{i}$ is available, a correction term must be introduced in the contrast function. Our approach, detailed in Section 3, involves replacing the observation of $X^{i}_{t_{j}}$ at each time $t_{j}$ with the increment of the rate process $\tilde{X}^{i}_{t_{j}}:=\frac{1}{\Delta_{n}}(Y^{i}_{t_{j+1}}-Y^{i}_{t_{j}})$, where $\Delta_{n}:=t_{j+1}-t_{j}$ for $j=0,\dots,n-1$ denotes the equidistant discretization step associated with a sample of size $n$.
@@
The outline of the paper is as follows. In Section 2, we introduce the model and outline the assumptions necessary for our main results, along with an explanation of how the propagation of chaos applies in this context. Section 3 presents our estimators and their properties, which constitute the core of our results. In Section 5, we discuss Hörmander’s condition and explain the transition from model (1) to (3), which is the model used for our estimation procedure. Section 6 introduces several key results that are required for proving our main findings, while Section 7 is dedicated to the proofs of these main results. Lastly, Appendix A focuses on the detailed proofs of all the technical results introduced throughout the paper.
@@ 2 Model and assumptions
We emphasize that, throughout our analysis, the time horizon $T$ will be fixed, and we will consider the asymptotic regime where only the number of particles $N$ and the sample size $n$ tend to infinity.
A key point in our analysis is that we observe discrete observations of the interacting particles, but we can transition from the interacting particle system described above to a system of independent particles through what is known as the propagation of chaos. The system of independent particles consists of $N$ i.i.d. copies of the McKean–Vlasov equation associated with the model (5), as detailed below.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{B}}(z,\pi):=\begin{pmatrix}x\\ b(z,\pi)\end{pmatrix}\in\mathbb{R}^{2},\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{A}}(z,\pi):=\begin{pmatrix}0&0\\ 0&a(z,\pi)\end{pmatrix}\in\mathbb{R}^{2\times 2}$
@@
We remark that the global Lipschitz assumption on the drift coefficient $b$ (Assumption A2) can be relaxed to accommodate locally Lipschitz drifts with polynomial growth. This relaxation is particularly relevant for interacting underdamped Langevin dynamics, where the potential function governing the drift is often non-linear, such as the standard double-well potential $V(x)=\frac{x^{4}}{4}-\frac{x^{2}}{2}$. In such a setting (cf. [28, Assumption 2.1]), the boundedness of moments established in Lemma 6.1 can be recovered by applying [29, Theorem 3.3], and the propagation of chaos follows from [28, Proposition 3.1]. Consequently, the main results of this paper remain valid under these weakened conditions.
Furthermore, throughout this work we focus on the one-dimensional kinetic setting where $Z_{t}^{i}=(X_{t}^{i},Y_{t}^{i})\in\mathbb{R}^{2}$, but the extension to the $d$-dimensional case ($Z_{t}^{i}\in\mathbb{R}^{2d}$) is theoretically consistent with our framework. Indeed, the propagation of chaos for such systems holds in any dimension, although the rate of convergence under the Wasserstein metric $W_{p}$ typically suffers from the curse of dimensionality, scaling as $N^{-1/d}$. However, since our derivation of the statistical rates depends on the convergence to the mean-field limit rather than the specific dimensional rate of that convergence, the results remain qualitatively unchanged. We have chosen to maintain the current one-dimensional notation to ensure the proofs remain accessible and to avoid the significant notational complexity associated with multidimensional matrix calculus.
We recall that the coefficients depend on two parameters, $\mu_{0}$ and $\sigma_{0}$, which are the objects of our estimation. Therefore, we will denote the drift coefficient as $b_{\mu_{0}}$ and the diffusion coefficient as $a_{\sigma_{0}}$. Furthermore, we will refer to the pair of parameters as $\theta_{0}:=(\mu_{0},\sigma_{0})$, which belongs to the interior of a set $\Theta:=\Theta_{1}\times\Theta_{2}$, where $\Theta_{i}\subset\mathbb{R}^{p_{i}}$, $i=1,2$, are compact and convex.
@@ 3 Estimators and main results
This brings us to the definition of the contrast function within the two observation frameworks under consideration. Recall that in the idealized setting where a continuous-time path of the process is assumed to be available, the maximum likelihood estimator provides a highly efficient framework for parameter estimation. However, in the realistic case where only a discrete-time sample of the process is observed, the transition density (and consequently the likelihood) is generally unavailable in closed form, making direct maximum likelihood estimation infeasible.
As pointed out in the introduction, a classical approach to overcoming this issue is to propose a quasi-likelihood function based on the Euler–Maruyama discretization of the process. Even in the case of complete observation, it is not feasible to directly apply the two-dimensional Euler contrast function for each particle $i$ to estimate the parameters $(\mu_{0},\sigma_{0})$. Specifically, the two-dimensional Euler–Maruyama approximation for any $i\in\{1,\dots,N\}$ is given by:
$\begin{pmatrix}Y^{i}_{(j+1)\Delta_{n}}\\ X^{i}_{(j+1)\Delta_{n}}\end{pmatrix}=\begin{pmatrix}Y^{i}_{j\Delta_{n}}\\ X^{i}_{j\Delta_{n}}\end{pmatrix}+\Delta_{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{B}}(Z^{i}_{j\Delta_{n}},\Pi^{N}_{j\Delta_{n}})+\sqrt{\Delta_{n}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{A}}(Z^{i}_{j\Delta_{n}},\Pi^{N}_{j\Delta_{n}})\begin{pmatrix}\xi^{i}_{j,1}\\ \xi^{i}_{j,2}\end{pmatrix}$,
@@
where $\theta=(\mu,\sigma)$ and the superscript $C$ indicates “complete observations”. This function is an extension of the classical Euler contrast for interacting particle systems in one dimension, as proposed in [4]. Starting from such contrast function, we define a minimum contrast estimator $\hat{\theta}_{n}^{N,C}$ for complete observations as
@@
where P stands for “partial observations”. Compared to the contrast function proposed for complete observations, the two key differences are the additional factor of $\frac{3}{2}$, which corrects for the dependence structure highlighted above (and further detailed in Remark 6.5), and a shift in the index of the drift term. It is important to note that $(\tilde{X}^{i}_{j\Delta_{n}})$ is not a Markov process, and the shift in the index of the drift coefficient has been introduced to avoid a correlation term between $(\tilde{X}^{i}_{(j+1)\Delta_{n}}-\tilde{X}^{i}_{j\Delta_{n}})$ and a function of $(\tilde{Z}^{i}_{j\Delta_{n}},\tilde{\Pi}^{N}_{j\Delta_{n}})$. Although this correlation term is small (of the order $\sqrt{\Delta_{n}}$), it is not negligible in our analysis, leading us to the contrast function as defined above.
@@
Assumptions A1A5 are necessary to establish the consistency of our estimator as detailed in Theorem 3.1. While these assumptions are standard in the statistical analysis of stochastic processes, the identifiability condition in Assumption A5 warrants particular attention. The quantities $I(\theta)$ and $J(\sigma)$, central to this requirement, depend on the underlying law of the system, making them challenging to verify explicitly in practice.
This challenge is common in parameter estimation for interacting systems; for a thorough discussion, we refer to [23, Section 2.4], where the authors link global identifiability to the non-degeneracy of the Fisher information matrix [23, Proposition 16]. To illustrate how these conditions can be checked, consider the example of an interacting underdamped Langevin dynamics with constant diffusion coefficient $a_{\sigma}(z,\pi)=\sqrt{2\gamma/\beta}$ and drift coefficient $b_{\mu}(z,\pi)=-\gamma x-V^{\prime}(y)-\int_{\mathbb{R}\times\mathbb{R}}W^{\prime}(y-v)\pi(du,dv)$, where the potentials $V(y)$, $W(y)$ are known and $\mu=\gamma$, $\sigma=\sqrt{\gamma/\beta}$. Then
$I(\theta)=\frac{(\mu_{0}-\mu)^{2}}{\sigma^{2}}\int_{0}^{T}\int_{\mathbb{R}\times\mathbb{R}}x^{2}\bar{\Pi}_{t}(dx,dy)dt\qquad\mbox{and }J(\sigma)=\left(\frac{\sigma_{0}^{2}}{\sigma^{2}}+\log\sigma^{2}\right)T.$
Under Assumption A3, $\inf_{\theta\in\Theta:|\mu-\mu_{0}|\geq\varepsilon}I(\theta)>0$ simply requires that the system does not remain at the origin, i.e. the second moment of the particles remains positive over the time horizon. Moreover, $J(\sigma)$ satisfies $\inf_{|\sigma-\sigma_{0}|\geq\varepsilon}(J(\sigma)-J(\sigma_{0}))>0$ for all $T>0$, since $\log x>1-\frac{1}{x}$ for all $x>0$, $x\neq 1$. Therefore, the identifiability conditions hold for this Langevin-type dynamics. Similarly, the conditions can be verified for the other examples in Section 4, where the drift and diffusion coefficients depend linearly on the unknown parameters.
@@
(Invertibility) The $p_{j}\times p_{j}$ matrices $\Sigma^{(j)}(\theta_{0})=(\Sigma^{(j)}_{kl}(\theta_{0}))$, defined by
$\Sigma^{(j)}_{kl}(\theta_{0}){:}=\begin{cases}\displaystyle\bar{\Pi}\Big(\frac{\partial_{\mu_{k}}b_{\mu_{0}}\partial_{\mu_{l}}b_{\mu_{0}}}{a^{2}_{\sigma_{0}}}\Big),\qquad&j=1,\,k,l=1,\dots,p_{1},\\ \displaystyle\bar{\Pi}\Big(\frac{\partial_{\sigma_{k}}a^{2}_{\sigma_{0}}\partial_{\sigma_{l}}a^{2}_{\sigma_{0}}}{a^{4}_{\sigma_{0}}}\Big),\qquad&j=2,\,k,l=1,\dots,p_{2},\end{cases}$
satisfy $\operatorname{det}(\Sigma^{{(j)}}(\theta_{0}))\neq 0$, $j=1,2$.
@@
$(\sqrt{N}(\hat{\mu}_{n}^{N,C}-\mu_{0}),\sqrt{N/{\Delta_{n}}}(\hat{\sigma}_{n}^{N,C}-\sigma_{0}))\xrightarrow{d}\mathcal{N}(0,(\Sigma^{(1)}(\theta_{0}))^{-1})\otimes\mathcal{N}(0,2(\Sigma^{(2)}(\theta_{0}))^{-1}).$
@@
$(\sqrt{N}(\hat{\mu}_{n}^{N,P}-\mu_{0}),\sqrt{N/{\Delta_{n}}}(\hat{\sigma}_{n}^{N,P}-\sigma_{0}))\xrightarrow{d}\mathcal{N}(0,(\Sigma^{(1)}(\theta_{0}))^{-1})\otimes\mathcal{N}(0,(9/4)(\Sigma^{(2)}(\theta_{0}))^{-1}).$
Here $\Sigma^{(j)}(\theta_{0})$, $j=1,2$, are as defined in A6.
@@
The proof of the main results is provided in Section 7 and centers on analyzing functionals of $Z^{i}_{t_{j}}=(Y_{t_{j}}^{i},X_{t_{j}}^{i})$ and $\tilde{Z}_{t_{j}}=(Y_{t_{j}}^{i},\tilde{X}_{t_{j}}^{i})$ for $i\in\{1,\ldots,N\}$ and $j\in\{1,\ldots,n-1\}$. The asymptotic analysis of these functionals proves to be challenging due to the involvement of both the Markovian process $Z^{i}_{t_{j}}$ and the non-Markovian process $\tilde{Z}^{i}_{t_{j}}$.
In our degenerate setting, the asymptotic variance involves the full law $\bar{\Pi}$ rather than merely the marginal of the second component. This mirrors the behavior of hypoelliptic diffusions in the ergodic $T\to\infty$ regime studied in [77], where the variance depends on the joint invariant density of $(X_{t},Y_{t})$. This contrasts with autonomous diffusions [56], where the asymptotic distribution depends only on the stationary measure of the observed component. This parallelism underscores how our findings for $N\to\infty$ align with the hypoelliptic structure identified in [77] (see Remark 3.4 for further details).
Our proposed drift estimator is efficient, as its asymptotic variance matches the Fisher information matrix derived from the Local Asymptotic Normality (LAN) property for $d$-dimensional McKean–Vlasov models under continuous observations [23]. Although the LAN property for joint estimation of drift and diffusion coefficients remains open, parallels with the classical $N=1$ and $T\to\infty$ setting (Remark 3.4) indicate that our diffusion estimator is inefficient. Specifically, constructing the contrast function solely from the rough coordinate inflates the asymptotic variance by a factor of $9/4$ relative to the optimal case. Thus, while consistent, the diffusion estimator lacks efficiency due to its partial use of the system’s information.